On February 1, 2021, Miter Corp. lends cash and accepts a $3,900 note receivable that offers 12% interest and is due in six months. How much interest revenue will Miter Corp. report during 2021? (Do not round intermediate calculations.)
Answer #1

Interest on notes receivable = $3,900 \times 12% \times 6/12 = $234

Hence interest revenue to reported during 2021 is $234